# Distributional Regression with Tabular Foundation Models: Evaluating Probabilistic Predictions via Proper Scoring Rules

### 09 Mar 2026 v1

@@ Abstract
Prior-Data Fitted Networks (PFNs), such as TabPFN and TabICL, have revolutionized tabular deep learning by leveraging in-context learning for tabular data. These models are meant as foundation models for classification and regression settings and promise to greatly simplify deployment in practical settings because their performance is unprecedented (in terms of mean squared error or $R^{2}$, when measured on common benchmarks like TabArena or TALENT).
However, we see an important weakness of current benchmarks for the regression setting: the current benchmarks focus on evaluating win rates and performance using metrics like (root) mean squared error or $R^{2}$. Therefore, these leaderboards (implicitly and explicitly) push researchers to optimize for machine learning pipelines which elicit a good mean value estimate.
The main problem is that this approach only evaluates a point estimate (namely the mean estimator which is the Bayes estimator associated with the mean squared error loss). In this article we discuss the application of proper scoring rules for evaluating the goodness of probabilistic forecasts in distributional regression. We also propose to enhance common machine learning benchmarks with metrics for probabilistic regression. To improve the status quo and make the machine learning community aware of scoring rules for probabilistic regression, we advocate to use the continuous ranked probability score (CRPS) in benchmarks for probabilistic regression. However, we also illustrate that the choice of the scoring rule changes the inductive bias of the trained model (see section 4.1). We, therefore, advocate for finetuning or promptable tabular foundation models.
@@ 1 Introduction
Besides the in-context learning, we recognize that both leading models TabPFN and also TabICLv2 perform forms of (discretized) distributional regression - an approach we greatly support since it allows for full probabilistic forecasts (in simple terms this means that the models allow to "predict a histogram over the support of potential outcomes", similar to the insets in figure 3).
@@
A great development of the emergence of TabPFN is that it moves away from point estimates — such as the conditional median (e.g. arising from mean absolute error minimization) or the conditional mean (from minimizing e.g. mean squared error)—to full distributional regression. Joining the previous rage against the mean, we agree that point estimates may not offer the deepest insight into a problemKneib et al. (2023) since they fail to capture aleatoric uncertainty (uncertainty you see from one-data point to another one).
We show a simple toy model to illustrate problems with the usual ordinary least squares estimation which tries to elicit a mean estimate: a regression problem where the target variable has multiple modes (see figure 3 and the appendix for the description of the data generating process and the models involved) [^2] One might argue that such a bimodal response may be due to an unobserved confounder. However, we still need to work with the data at hand and, therefore, we argue that modeling the aleatoric uncertainty is important to make better decisions..
@@
Figure 3: The images illustrate effects of training with CRPS or log-score (cross entropy loss) with same training data, same architecture but other loss function (CRPS vs log score) at finite amount of samples (1600 train samples, 400 test samples). For further description of the models and the data generating process see appendix.
@@
One main takeaway from the toy model in figure 3 is that mean value estimates may fall in regions without realizations and may (depending on the utility) be completely useless (think about asking: "What could be the next outcome of a six sided die?" Even if you had a perfect mean value estimate telling you the mean value 3.5, then this would not even be a valid realization). Additionally, we also see that the usual construction of a prediction interval around the mean value estimate would cover a lot of empty space without samples (just by construction as an interval), which is inferior to estimating a conditional histogram like shown in the figure.
@@
Interestingly, for the toy dataset, the Cross-entropy loss trained model in figure 3 differs significantly in terms of alignment with the ground truth data generating process, which indicates that at finite amount of samples:
@@
  • the convergence speed during model training may depend on the chosen loss function - even if both loss functions derive from proper scoring rules
@@
  • the sample efficiency of the complete training pipeline may depend on the loss function - even if both loss functions derive from proper scoring rules
@@
In the table 11 in the appendix we observe that different models have different performance metrics, depending on the loss function even at otherwise identical training procedure. We also had observed this behavior on other openml datasets [^3] See https://github.com/PriorLabs/TabPFN/pull/689 for the pull request proposing CRPS for fine-tuning TabPFN.
However, in practice, we minimize the sample average score $\widehat{E[\mathbf{S}]}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}(F,y_{i})$. While $\widehat{E[\mathbf{S}]}$ is a consistent estimator of the expectation, the optimization path at finite $n$ is governed by the specific gradients $\nabla\mathbf{S}$ induced by the rule.
The Logarithmic Score, $\mathbf{S}_{Log}(F,y)=-\log f(y)$, exhibits extreme sensitivity to the tails of $\Omega$. Since $\nabla\mathbf{S}_{Log}\propto 1/f(y)$, the penalty for low-density observations grows toward infinity. At finite $n$, even a single outlier can dominate the gradient, forcing the optimizer to prioritize tail-fitting early, often at the expense of the distribution’s central mass.
Conversely, the Continuous Ranked Probability Score (CRPS) provides a more robust optimization landscape. This is best understood by expressing the CRPS as an integral over pinball losses $\rho_{\tau}$ for all quantiles $\tau\in[0,1]$:
$\text{CRPS}(F,y)=2\int_{0}^{1}\rho_{\tau}(F^{-1}(\tau),y)d\tau,$ (2)
where $\rho_{\tau}(u,y)=(y-u)(\tau-\mathbb{I}\{y<u\})$.
Because the gradient of the pinball loss is bounded (constant in magnitude), the CRPS gradient remains stable regardless of how far $y$ lies in the tails. Consequently, at finite $n$, CRPS-based estimation focuses on the global probability mass and is less susceptible to being pulled away by sparse observations in $\Omega$. While both rules are asymptotically equivalent in their target, their finite-sample "optimum" differs and also their behavior for tail-sensitivity (LogS) and robust mass-estimation (CRPS).
1.1 Important Quotes
  • "It is difficult to understand why statisticians commonly limit their inquiries to averages, and do not revel in more comprehensive views. Their souls seem as dull to the charm of variety as that of the native of one of our flat English counties, whose retrospect of Switzerland was that, if its mountains could be thrown into its lakes, two nuisances would be got rid of at once." (Galton, 1889) Kneib et al. (2023)
  • "Despite these well-argued perspectives, there has been little recognition that the common practice of requesting some point forecast, and then evaluating the forecasters by using some (set of) scoring function(s), is not a meaningful endeavor. In this article, we develop the perspectives of Murphy and Daan (1985) and Engelberg, Manski, and Williams (2009) and argue that effective point forecasting depends on guidance or directives, which can be given in one of two complementary ways, namely, by disclosing the scoring function ex ante to the forecaster, or by requesting a specific functional of the forecaster’s predictive distribution, such as the mean or a quantile." Gneiting (2011)
  • "Although every proper scoring rule can identify the true model or the conditional distribution as its minimum in the population sense, they exhibit different inductive biases or regularization effects in finite samples. Furthermore, the computational tractability varies substantially depending on the chosen scoring rule" Waghmare and Ziegel (2025)
1.2 Contributions
We greatly appreciate the recent improvement of tabular foundation models to move away from point estimates.
However, current regression benchmarks for Tabular Machine Learning like TabArena or TALENT are still stuck in evaluating "historically important" point estimates only.
We
  • propose the enhancement of regression benchmarks to account for probabilistic regression metrics derived from scoring rules, mainly CRPS by discussing a bimodal toy model.
  • we provide first evaluations of realTabPFNv2.5 and TabICLv2 in terms of proper scoring rules (CRPS and beta energy score).
  • we show how finetuning with different scoring rules than the one seen during pretraining in a finite-sample setting can enhance realTabPFNv2.5 performance (CRLS, beta energy score)
  • we show that in theory the choice of a scoring rule can have a significant impact in terms of the inductive bias of the model (having implications for finetuning and benchmarking)Merkle and Steyvers (2013); Waghmare and Ziegel (2025)
  • Additionally in the outlook, we start the discussion around a potential conflict between optimality of predictions with regard to user-supplied scoring rule and the goal to not adapt a general-purpose foundation model after pretraining.
@@ 2.1 Log Score
Viewing discretized regression as a multi-class classification problem suggests the application of Cross-Entropy (CE), which corresponds to the log scoring rule which is local. This choice may be optimal if the final business objective is strictly to maximize the probability of the "correct" bin.
@@ 2.2 Continuous Ranked Probability Score (CRPS)
However, using cross entropy loss (meaning the log scoring rule) would ignore the distance between the prediction and the ground truth (bins are not even treated as having an ordinal order like in $1<2<3$ but more like completely independent concepts like "panda", "tiger", "fish"). To inject geometric awareness into the gradient, we proposed to use CRPS for finetuning TabPFN in December 2025444https://github.com/PriorLabs/TabPFN/pull/689 and saw CRPS being used in TabICLv2 training in February 2026Qu et al. (2026). In a regression context, predicting $y=100$ when the truth is $y=10$ should be penalized more heavily than predicting $y=11$. Still, standard Cross-Entropy treats both errors similarly if they fall into different wrong bins.
@@
where $F(x_{i})$ is the predicted cumulative distribution function (e.g. obtained via cumsum of the softmax output of a neural network over the binned support). To perform optimum score estimation the expectation value (over all training samples) scoring rule needs to be taken and minimized.
However, CRPS weights all quantiles of the output space equally. This is likely, prohibitive in applications with asymmetric risk structure over the output space (see later discussion also for weighted CRPSGneiting and Ranjan (2011)).
@@ 2.3 Continuous Ranked Logarithmic Scoring Rule (CRLS)
CRLS is defined by [^5] https://scoringrules.readthedocs.io/en/latest/theory.html
$\mathrm{CRLS}(F,y)=-\int_{-\infty}^{\infty}\log|F(x)+\mathbf{1}\{y\leq x\}-1|dx.$ (4)
@@ 2.4 Multivariate Enhancements
For $y\in\Omega\subseteq\mathbb{R}^{d}$ with $d>1$, univariate scores are insufficient for capturing dependence. We employ and suggest to benchmark for the Energy Score ($\mathbf{S}_{\beta}$) and Variogram Score ($\mathbf{S}_{VS,p}$) as also done in Landsgesell (2026).

### 13 Mar 2026 v2

@@ Abstract
Tabular foundation models such as TabPFN and TabICL already produce full predictive distributions, yet the benchmarks used to evaluate them (TabArena, TALENT, and others) still rely almost exclusively on point-estimate metrics (RMSE, $R^{2}$). This mismatch implicitly rewards models that elicit a good conditional mean while ignoring the quality of the predicted distribution.
We make two contributions. First, we propose supplementing standard point metrics with proper scoring rules (CRPS, CRLS, and the Interval Score) and provide a head-to-head comparison of realTabPFNv2.5 and TabICLv2 with regards to some proper scoring rules across 20 OpenML regression datasets. Second, we show analytically and empirically that different proper scoring rules induce different model rankings and different inductive biases during training, even though each rule is individually minimized by the true distribution. Fine-tuning realTabPFNv2.5 with scoring rules not seen during pretraining (CRLS, $\beta{=}1.8$ energy score) yields consistent improvements on the corresponding metrics, confirming that the training loss shapes the model beyond what propriety alone guarantees. Together, these findings argue for (i) reporting distributional metrics in tabular regression benchmarks and (ii) making the training objective of foundation models adaptable (via fine-tuning or task-token conditioning) to the scoring rule relevant to the downstream decision problem.
@@ 1 Introduction
Both leading models, TabPFN and TabICLv2, perform forms of discretized distributional regression, an approach we strongly support, as it enables full probabilistic forecasts (i.e., the models can predict a histogram over the support of potential outcomes, as illustrated by the insets in Figure 3).
@@
A notable contribution of TabPFN is its shift away from point estimates (such as the conditional median arising from MAE minimization, or the conditional mean from MSE minimization) toward full distributional regression. We agree that point estimates often fail to offer the deepest insight into a problem Kneib et al. (2023), as they cannot capture aleatoric uncertainty, i.e., the inherent variability from one data point to another. As Galton already noted in 1889:
“It is difficult to understand why statisticians commonly limit their inquiries to averages, and do not revel in more comprehensive views. Their souls seem as dull to the charm of variety as that of the native of one of our flat English counties, whose retrospect of Switzerland was that, if its mountains could be thrown into its lakes, two nuisances would be got rid of at once.” Kneib et al. (2023)
Recent work has explored a range of strategies for adapting tabular foundation models, including meta-learning, supervised fine-tuning, and parameter-efficient fine-tuning (PEFT) Tanna et al. (2026, 2025). However, to our knowledge, these studies focus on whether and how to fine-tune (in terms of parameter update strategies), while the inductive bias introduced by the choice of loss function (scoring rule) for a probabilistic forecast has not been systematically investigated.
1.1 Contributions
Current regression benchmarks for tabular machine learning (e.g. TabArena, TALENT) evaluate models exclusively through point-estimate metrics such as RMSE or $R^{2}$. This practice implicitly rewards models that elicit a good conditional mean while ignoring the quality of the full predictive distribution, a critical shortcoming given that leading tabular foundation models (TabPFN, TabICLv2) already produce distributional outputs.
Our contributions are twofold:
(I) Towards probabilistic regression benchmarks.
  • We argue, through a bimodal toy model and 20 OpenML regression datasets, that point-estimate leaderboards are insufficient and can be misleading when the predictive distribution is multimodal or heavy-tailed.
  • We propose a concrete benchmark enhancement that supplements MAE, RMSE, and $R^{2}$ with distributional metrics (CRPS, CRLS, and Interval Score, $\mathrm{IS}_{95}$) and provide a head-to-head evaluation of realTabPFNv2.5 and TabICLv2 using proper scoring rules.
  • We additionally outline multivariate extensions (Energy Score, Variogram Score) for benchmarking future joint predictive distributions.
(II) Scoring-rule inductive bias and the need for adaptable training objectives.
  • We show analytically (via the $\beta$-energy score family) and empirically that different proper scoring rules induce different model rankings, even though each rule is individually minimized by the true distribution Merkle and Steyvers (2013); Waghmare and Ziegel (2025). This may imply that no single pretraining loss can be universally optimal across all downstream evaluation criteria - it is essentially the question of how different errors are penalized and a global model may not be optimal for every application.
  • We demonstrate that fine-tuning realTabPFNv2.5 with scoring rules not seen during pretraining (CRLS, $\beta{=}1.8$ energy score) yields consistent improvements over the base model on the corresponding metrics across the majority of the 20 benchmark datasets, confirming that the inductive bias of the training loss can be shifted post hoc.
2 Motivating Example: Why Point Estimates Fail
We illustrate the limitations of point-estimate evaluation with a simple toy model: a regression problem where the target variable has multiple modes (see Figure 3 and the appendix for the description of the data generating process and the models involved) [^2] One might argue that such a bimodal response may be due to an unobserved confounder. However, we still need to work with the data at hand and, therefore, we argue that modeling the aleatoric uncertainty is important to make better decisions..
@@ 2 Motivating Example: Why Point Estimates Fail
Figure 3: The images illustrate the effects of training with CRPS or log score (cross-entropy loss) with the same training data and architecture but a different loss function (CRPS vs. log score) at a finite number of samples (1600 train samples, 400 test samples). For further description of the models and the data generating process, see the appendix.
@@
One main takeaway from the toy model in Figure 3 is that mean value estimates may fall in regions without realizations and may (depending on the utility) be completely useless (think about asking: “What could be the next outcome of a six-sided die?” Even if you had a perfect mean value estimate telling you the mean value 3.5, then this would not even be a valid realization). Additionally, the usual construction of a prediction interval around the mean value estimate would cover a lot of empty space without samples (just by construction as an interval), which is inferior to estimating a conditional histogram like shown in the figure.
@@
Interestingly, for the toy dataset, the cross-entropy loss trained model in Figure 3 differs significantly in terms of alignment with the ground truth data generating process, which indicates that at a finite number of samples:
@@
  • the convergence speed during model training may depend on the chosen loss function, even if both loss functions derive from proper scoring rules
@@
  • the sample efficiency of the complete training pipeline may depend on the loss function, even if both loss functions derive from proper scoring rules
@@
In Table 5 in the appendix we observe that different models have different performance metrics depending on the loss function, even under otherwise identical training procedures. This behavior has also been observed on other OpenML datasets [^3] See https://github.com/PriorLabs/TabPFN/pull/689 for the pull request proposing CRPS for fine-tuning TabPFN..
@@ 3.1 Log Score
Viewing discretized regression as a multi-class classification problem suggests the application of cross-entropy (CE), which corresponds to the log scoring rule, which is local. This choice may be optimal if the final business objective is strictly to maximize the probability of the "correct" bin.
@@ 3.2 Continuous Ranked Probability Score (CRPS)
However, using cross-entropy loss (meaning the log scoring rule) would ignore the distance between the prediction and the ground truth (bins are not even treated as having an ordinal order like in $1<2<3$ but more like completely independent concepts like "panda", "tiger", "fish"). To inject geometric awareness into the gradient, we proposed to use CRPS for fine-tuning TabPFN in December 2025444https://github.com/PriorLabs/TabPFN/pull/689 and saw CRPS being used in TabICLv2 training in February 2026 Qu et al. (2026). In a regression context, predicting $y=100$ when the truth is $y=10$ should be penalized more heavily than predicting $y=11$. Still, standard cross-entropy treats both errors similarly if they fall into different wrong bins.
@@
where $F(x_{i})$ is the predicted cumulative distribution function (e.g. obtained via cumsum of the softmax output of a neural network over the binned support). To perform optimum score estimation, one minimizes the empirical mean of the scoring rule over all training samples.
However, CRPS weights all quantiles equally, which can be limiting in applications with asymmetric risk over the output space (see the discussion of weighted CRPS below Gneiting and Ranjan (2011)).
@@
3.3 Weighted CRPS (wCRPS)
Accounting for asymmetric risk structures, a modeler can choose to apply the weighted CRPS Gneiting and Ranjan (2011), a family of strictly proper scoring rules that enables emphasizing prediction quality for the left or right tail of a distribution. This makes sense when dealing with e.g. energy forecasting or forecasting weather events with asymmetric cost. In section 4.2 we discuss the practical implications of the choice of the scoring rule for the inductive bias of the model and for fine-tuning. However, here we only define it as a potential alternative to the CRPS for fine-tuning and benchmarking (while not using it since weighting would be use-case dependent):
@@ 3.4 Continuous Ranked Logarithmic Scoring Rule (CRLS)
CRLS (also called "Exceedance Probability Score", EPS) was introduced by Juutilainen et al. (2012) [^5] https://scoringrules.readthedocs.io/en/latest/theory.html
$\mathrm{CRLS}(F,y)=-\int_{-\infty}^{\infty}\log|F(x)+\mathbf{1}\{y\leq x\}-1|dx,$ (3)
where a negation is introduced relative to the original paper so that lower scores correspond to better forecasts (minimization convention).
"The exceedance probability score differs from CRPS so that it evaluates the difference between the binary event and its predicted probability using the logarithm rather than the negative squared difference." Juutilainen et al. (2012) "[CRLS] gives more weight to events [than CRPS] that were predicted to be very unlikely. [CRLS] differs significantly also from the log scoring rule: [CRLS] is distance sensitive, whereas [Log score] is not." Juutilainen et al. (2012) "CRPS is very insensitive to an error in which predicted deviation is too low and that [Log score] is relatively insensitive to an error in which predicted deviation is too high. [CRLS] penalizes more on crude errors in predicted mean than LogS or CRPS." Juutilainen et al. (2012)
3.5 Interval Score
The Interval Score Gneiting and Raftery (2007) targets the $(1-\alpha)$ central prediction interval and is strictly proper, directly answering “how well is the interval estimated?” for a given significance level $\alpha$:
where $[l,u]$ are the $\alpha/2$ and $1-\alpha/2$ quantiles of the prediction $F$. We report $\mathrm{IS}_{95}$ (i.e. $\alpha=0.05$) as one of our six evaluation metrics.
This scoring rule highlights that different scoring rules favor optimization in distinct regions of the PMF; in the extreme case of the interval score, only the two boundary quantiles are penalized. General-purpose models may therefore fail to achieve optimality for tasks where the cost of error is non-uniform across $\Omega$.
@@ 3.6 Multivariate Enhancements
For $y\in\Omega\subseteq\mathbb{R}^{d}$ with $d>1$, univariate scores are insufficient for capturing dependence. We employ and suggest benchmarking the Energy Score ($\mathbf{S}_{\beta}$) and Variogram Score ($\mathbf{S}_{VS,p}$), as also done in Landsgesell (2026).
@@
4 Choice of Scoring Rules for Decision Making
Different proper scoring rules penalize different types of errors differently and thus induce different inductive biases Waghmare and Ziegel (2025); Merkle and Steyvers (2013). As Gneiting emphasizes:
“Despite these well-argued perspectives, there has been little recognition that the common practice of requesting some point forecast, and then evaluating the forecasters by using some (set of) scoring function(s), is not a meaningful endeavor. […] effective point forecasting depends on guidance or directives, which can be given in one of two complementary ways, namely, by disclosing the scoring function ex ante to the forecaster, or by requesting a specific functional of the forecaster’s predictive distribution, such as the mean or a quantile.” Gneiting (2011)
Several lines of work address how to choose the right rule: Oesterheld Oesterheld and Conitzer (2020) formalizes “Decision Scoring Rules”; Johnstone et al. Johnstone et al. (2011) derive tailored scoring rules from a utility perspective (their orchard weather-forecast example on page 2 is particularly instructive); and Merkle et al. Merkle and Steyvers (2013) show empirically that the ranking of forecasters changes with the scoring rule (the Spearman correlation between Brier-score and log-score rankings in their Table 1 is only 0.15).
We port this idea to the regression setting using the $\beta$-energy score family and show that the optimal model depends on the choice of $\beta$.
4.1 Finite-Sample Behavior of Proper Scoring Rules
In practice, we minimize the sample average $\widehat{E[\mathbf{S}]}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}(F,y_{i})$, whose optimization path at finite $n$ depends on the gradients $\nabla\mathbf{S}$.
Two properties distinguish the rules used in this paper:
  • The Logarithmic Score, $\mathbf{S}_{Log}(F,y)=-\log f(y)$, has gradient $\nabla\mathbf{S}_{Log}\propto 1/f(y)$, which is unbounded as $f(y)\to 0$.
  • The CRPS, expressed as $\text{CRPS}(F,y)=2\int_{0}^{1}\rho_{\tau}(F^{-1}(\tau),y)\,d\tau$ with pinball loss $\rho_{\tau}(u,y)=(y-u)(\tau-\mathbb{I}\{y<u\})$, has a gradient that is bounded in magnitude.
The unbounded gradient of the log score means that low-density observations contribute disproportionately large updates, whereas CRPS weights all quantile levels equally. Both rules are strictly proper and thus share the same population-level minimizer, but their finite-sample optimization dynamics differ due to these gradient properties.
4.2 Investigating the effect of the choice of scoring rule on the optimal forecast
$\hat{p}_{i}(Y=y|X=x)=\delta(y-f_{i}(x)),$ (8)
$p_{g}(Y=y|X=x)=\delta(y-g(x)).$ (9)
We choose $x\in[-3,3]$777x=np.linspace(-3,3,1000).
  • $g(x)=\sin(x)$
  • Model A: Small consistent bias, $f_{A}(x)=g(x)+0.1$
  • Model C: Scaled version of the true function, $f_{C}(x)=0.8\cdot g(x)$
  • Model D: Shifted version of the true function, $f_{D}(x)=\sin(x+0.2)$
  • Model E: Simple baseline at zero, $f_{E}(x)=0$
A practitioner would naturally ask which model is best. We examine the models in Figure 4 for intuition.
The distinct error profiles of the models are visible in Figure 5. Some models make small errors everywhere (model A), some models make large errors in a small region (model B), some models make medium-sized errors everywhere (model C) and some models make small errors everywhere but are shifted (model D).
Figure 5: Absolute errors for the toy models
Now we can apply different scoring rules (from the $\beta$-energy family) to see which model is optimal for which scoring rule. For a fixed choice of $\beta$ (one point on the x-axis in Figure 6) the better model has the lower beta-energy score.
Figure 6: Decision based on scoring rules for the toy models
We observe that Model B ranks best for $\beta\lesssim 1$, but for $\beta\gtrsim 1$ the ranking changes, with other models also switching rank (Figure 7).
Figure 7: Ranking of toy models based on scoring rules
Fundamentally this behavior arises from different scoring rules rating the severity of errors in Figure 5 differently (and there are even scoring rules which penalize errors depending on the direction of the error, see for example asymmetric CRPS Iacopini et al. (2023) which may be important e.g. in finance where errors have different severity depending on the direction of the error: making 100 percent more money is less severe than losing 100 percent of money).
For simple probabilistic point masses like shown here we can even make an analytical statement: The beta-energy score with $\beta=1$ favors minimization of the mean absolute error (i.e. prefers estimating a median estimate) while for $\beta=2$ the scoring rule prefers a minimization of mean squared error.
In Table 1 we also note explicitly the ranking of models changes for three selected $\beta$ values:
Table 1: Model Rankings for Different $\beta$ Values
Model Rank ($\beta=0.2$) Rank ($\beta\approx 1.0$) Rank ($\beta=2.0$)
Model A (Bias) 2 1 1
Model B (Outlier) 1 4 4
Model C (Scale) 4 3 3
Model D (Shift) 3 2 2
Model E (Zero) 5 5 5
Derivation.
where $X$ and $X^{\prime}$ are independent draws from $P$.
$\mathrm{ES}_{\beta}(\delta_{\hat{m}},\,y)=|\hat{m}-y|^{\beta}.$ (11)
Summary.
The choice of scoring rule changes the ranking of forecasting models: in this toy model the result follows analytically from the $\beta$-energy score reducing to an $L^{\beta}$ norm. Whether this ranking sensitivity carries over to trained distributional models in realistic tabular settings remains an open empirical question, but both the theoretical argument and the empirical literature Waghmare and Ziegel (2025); Merkle and Steyvers (2013) support the view that the choice of scoring rule for pretraining or fine-tuning a tabular foundation model is non-trivial and may affect which use-cases the model is best suited for Oesterheld and Conitzer (2020); Johnstone et al. (2011). As Waghmare et al. summarize:
“Although every proper scoring rule can identify the true model or the conditional distribution as its minimum in the population sense, they exhibit different inductive biases or regularization effects in finite samples. Furthermore, the computational tractability varies substantially depending on the chosen scoring rule.” Waghmare and Ziegel (2025)
How to best choose a scoring rule for general-purpose tabular foundation model pretraining and how to enable adaptation to business cases with asymmetric error severity remains an open question; we discuss possible routes in section 6.